Car A left Melbourne for Adelaide at 11.00 am and traveled at at an average speed of 70 km/h.
Car B left Melbourne for aDELAIDE at 1.00 pm same day and traveled at an average speed of 90 km/h.
At what time will Car B catch Car A?

Answer by mananth(16949)

Car A 70 mph
Car B 90 mph
Car A 11:00
Car B 01:00
Difference in time= 14:00 => 2.00 hours
Car A will have covered 140.00 miles when Car B starts
catch up distance= 140.00 miles
catch up speed = 90 -70
catch up speed = 20 mph
Catchup time = catchup distance/catch up speed
catch up time= 140/20= 7
catch up time= 7.00 hours
They will meet at 08:00 pm